OpenStudy (anonymous):

1/2 + 2/x = 3/4, the answer is 8 but how do you get it?

hero (hero):

\[\frac{1}{2} + \frac{2}{x} = \frac{3}{4}\] Subtract 1/2 from both sides: \[\frac{2}{x} = \frac{3}{4} - \frac{1}{2}\] 1/2 = 2/4 \[\frac{2}{x} = \frac{3}{4} - \frac{2}{4}\] \[\frac{2}{x} = \frac{1}{4}\] Cross Multiply: \[(2)(4) = 1(x)\]\(8 = x\)

OpenStudy (anonymous):

thank you

hero (hero):

yw